# Large Language Model-Assisted UAV Operations and Communications: A Multifaceted Survey and Tutorial

### v1

@@ III-D Multi-LLM Architectures for UAV Communications
Multi-LLM systems are architectures in which multiple LLMs collaborate, either in parallel or sequentially, to solve complex tasks. Unlike single-model approaches, multi-LLM systems leverage the complementary strengths of individual models, enabling specialization, scalability, robustness, and effective task decomposition. Performance is further improved through cross-verification, which helps detect errors, reduce bias, and mitigate hallucinations. Collaboration among LLMs can follow cooperative, competitive, or ensemble paradigms, while system architectures may be centralized or decentralized, depending on fault tolerance and task allocation requirements. Additionally, multi-LLM frameworks can enhance privacy by routing sensitive queries to trusted models while validating outputs across agents.
@@
Extending multi-LLM coordination to interconnected aerial networks, Yan et al. [126] propose a hierarchical control framework for hybrid terrestrial–non-terrestrial systems incorporating High-Altitude Platform Stations (HAPS). Specifically, a meta-controller LLM deployed on the HAPS manages global resource allocation and load balancing, while onboard LLMs handle real-time motion planning and communication decisions. Experimental results show that this collaborative architecture achieves higher system rewards, lower operational costs, and lower collision rates than DRL-assisted and hybrid baselines. Fig. 6 depicts Multi-LLM architectures for UAV communications, and Table IX provides a summary of the works.
@@
Table IX: Hierarchical and multi-LLM architectures for coordinated UAV and aerial network operations, comparing distributed, swarm-based, and layered LLM frameworks in inspection, security, and hybrid terrestrial–non-terrestrial network applications.
@@ III-E Key Findings and Insights
III-E1 Integrated Communication–Perception–Reasoning Architectures for Intelligent UAV Communications
Recent UAV platforms integrating 5G connectivity, edge AI, immersive interfaces, and LLM-based reasoning demonstrate a fundamental shift from modular system designs toward tightly coupled communication–perception–reasoning architectures. Rather than treating communication, sensing, and intelligence as independent components, these systems show that their close integration is essential for achieving real-time situational awareness, closed-loop human–UAV interaction, and adaptive decision support in latency-sensitive and bandwidth-intensive environments. This observation indicates that communication must be designed as a first-class enabler of interactive and intelligent aerial systems, rather than a standalone data-delivery service.
III-E2 LLM-assisted Spectrum Management
@@
Recent frameworks for LLM-assisted green spectrum management in 6G Non-Terrestrial Networks (NTNs) integrate LLMs into a hierarchical three-layer architecture comprising data collection, intelligent LLM processing, and feedback-enabled execution. This design enables dynamic and adaptive spectrum allocation across satellites, high-altitude platforms, UAVs, and terrestrial networks. A key innovation is the joint optimization of spectral efficiency and energy consumption, supported by mechanisms including real-time interference analysis, predictive energy modeling, and an explicit energy–spectrum trade-off function. Simulation results show that the LLM-assisted approach significantly outperforms conventional methods, supporting up to 175 active transmitters – a 167% increase in capacity – while maintaining stable energy consumption and tighter interference control under varying threshold conditions. These findings advance intelligent and sustainable spectrum management for future 6G NTNs and underscore the transformative potential of LLMs in enabling green, scalable, high-performance global communications [97].
III-E3 LLM-assisted Near-Field Communications
@@
The integration of LLMs into near-field communications for the Low-Altitude Economy (LAE) brings transformative capabilities that address the unique challenges of UAV-centric wireless networks. By analyzing high-dimensional channel-state information in real time, LLMs can intelligently classify users as near-field or far-field without requiring precise distance estimation, thereby reducing computational overhead. They also enhance multi-user precoding by learning to predict beamforming and power allocation parameters directly from channel data, enabling efficient real-time adaptation to dynamic UAV mobility. Leveraging strong feature-extraction and pattern-recognition capabilities, LLMs can process the massive amounts of antenna and user data generated by extremely large-scale Multiple-Input Multiple-Output (MIMO) systems, thereby enabling performance to scale effectively as network density increases. In addition, their adaptability and few-shot learning abilities enable generalization across diverse operating conditions, such as varying UAV densities or interference levels, without extensive retraining. Finally, LLMs support unified multi-task learning by jointly handling user classification, precoding, and interference management within a single model, significantly simplifying overall system design [124].
III-E4 LLM-assisted ISAC Optimization
@@
Overall, these studies highlight a paradigm shift in ISAC optimization, moving beyond purely model-based or data-driven techniques. Instead of serving as standalone solvers, LLMs function as powerful complementary components that enhance traditional optimization frameworks. Their main strength is in addressing problem segments with combinatorial complexity, non-convexity, or a lack of explicit gradient information, where implicit trade-off reasoning is crucial. By encoding domain knowledge and constraints via natural-language prompting, LLMs provide a flexible reasoning interface that reduces reliance on extensive mathematical reformulation and scenario-specific expert tuning.
III-E5 High-Speed Connectivity for LLM-assisted UAVs
@@
Nevertheless, despite the advantages of high-speed connectivity, UAV-assisted LLM systems still face intrinsic challenges when handling long data sequences. Transformer-based LLMs have quadratic computational and memory complexity with respect to input sequence length, making long-context processing increasingly expensive even when inference is partially offloaded. As a result, extended missions, continuous perception streams, and prolonged human–UAV interactions remain difficult to support efficiently. Therefore, addressing long-context modeling and efficient attention mechanisms is essential to fully exploit LLMs’ real-time, 5G-assisted capabilities in dynamic UAV operations.
III-E6 Assessing LLM Reasoning and Safety
Field deployments of integrated LLM-enabled UAV systems confirm significant improvements in operational adaptability and effectiveness across diverse applications, including emergency response, environmental monitoring, and infrastructure inspection. At the same time, these evaluations expose persistent constraints related to bandwidth availability, energy consumption, multimodal robustness, and reasoning reliability. These observations highlight the necessity of holistic co-design across communication, computation, and control, as well as the importance of standardized evaluation frameworks, such as UAVBench and UAVThreatBench, to objectively assess reasoning performance and cybersecurity readiness for safe and scalable deployment of intelligent UAV communication systems.
UAVBench marks a significant advancement for the autonomous aerial systems community by providing the first large-scale, standardized, and physically grounded framework for evaluating LLM reasoning in realistic UAV mission scenarios. Prior to UAVBench, no unified methodology existed to assess whether LLMs genuinely understood the critical constraints governing aerial autonomy, such as aerodynamics, flight safety, navigation limits, airspace structure, and ethical decision-making. This absence of standardized evaluation made it difficult to distinguish models capable of supporting trustworthy UAV operations from those that generate superficially plausible but unreliable responses.
The benchmark introduced in [34] addresses this gap by providing 50,000 validated UAV scenarios and an additional 50,000 multiple-choice questions ($UAVBench_{MCQ}$) spanning 10 distinct reasoning styles. This comprehensive design enables systematic evaluation of a wide range of cognitive capabilities, from basic situational awareness to advanced multi-agent coordination and ethics-aware decision-making under operational constraints. Empirical evaluation of 32 state-of-the-art LLMs reveals that, while current models perform relatively well on perception-related and policy-level reasoning tasks, they continue to struggle with ethical judgment, resource-constrained planning, and the complex trade-offs inherent in real-world aerial missions. By releasing both the benchmark and its supporting tools, UAVBench enables the community to diagnose model limitations, compare approaches on a standard foundation, and guide the development of next-generation UAV intelligence through rigorous, transparent evaluation.
Complementing UAVBench, UAVThreatBench introduces a structured benchmark for evaluating LLM performance in UAV cybersecurity within industrial operational technology environments, in accordance with the European Radio Equipment Directive (RED). The benchmark comprises 924 expert-curated scenarios, each annotated with five distinct threats, yielding 4,620 RED-mapped threats across network and device integrity, privacy protection, and fraud prevention. Seven state-of-the-art LLMs, including OpenAI’s GPT-4 variants and LLaMA-3 models, were evaluated on 100 scenarios using fuzzy matching to assess alignment with expert-defined threats. Instruction-tuned models, such as GPT-4o, achieved the highest performance, reaching nearly 87% total threat matching, while open-weight models showed lower alignment and particular difficulty with availability and backend-layer threats. Further semantic analysis revealed systematic gaps in detecting denial-of-service vectors and vulnerabilities in MES/WMS logic.
Together, UAVBench and UAVThreatBench provide complementary, standardized evaluation frameworks that address both operational reasoning and cybersecurity resilience. By offering objective, mission-relevant benchmarks, these contributions equip the UAV research community with tools to assess, compare, and improve LLM-assisted autonomy, paving the way for more reliable, responsible, and deployment-ready intelligent aerial systems.
@@ IV MLLMs: Integrating Vision, Language, and Reasoning in UAV Operations and Communications
Recent advances in MLLMs have demonstrated that jointly reasoning over language, vision, and heterogeneous sensor inputs is critical for aerial platforms operating in dynamic, partially observable environments. In UAV applications, this multimodal integration enables intuitive human–machine interaction, robust perception and navigation, and adaptive mission execution by grounding abstract instructions in real-time sensory observations. Compared with text-only models, MLLMs provide a more natural interface between human intent and autonomous aerial behavior, particularly in scenarios involving VLN, swarm coordination, and mission-level planning.
Building on this foundation, this section reviews the core architectural principles of MLLMs, practical training and adaptation strategies for aerial data, and emerging multimodal reasoning mechanisms such as Multimodal Chain-of-Thought (MCoT) and Multimodal In-Context Learning (M-ICL). We then survey representative UAV applications enabled by MLLMs, including human–swarm interaction, VLN, autonomous mission planning, and real-time operational intelligence. Finally, we discuss key challenges related to safety, efficiency, and reliability that must be addressed to enable robust deployment of multimodal intelligence on resource-constrained aerial platforms.
@@ IV-A Building and Fine-Tuning Multimodal LLMs
A key challenge lies in aligning heterogeneous modalities for grounded reasoning. Contemporary MLLMs typically rely on modality-specific encoders coupled with a pretrained LLM, enabling language modules to attend selectively to visual features. This flexible architecture serves as a common backbone for UAV perception and reasoning tasks, but its effectiveness in aerial settings depends strongly on downstream adaptation.
Following pretraining, models are fine-tuned on UAV-specific tasks such as aerial image captioning, visual question answering, and object recognition. Dedicated datasets including AeroCaps [8] and HRVQA-VL [38] support fine-grained scene understanding and multi-hop reasoning over UAV imagery. Empirical evidence shows that task-aware fine-tuning is critical: compact MLLMs (e.g., 2B parameters) adapted with UAV data can outperform significantly larger generic models while requiring only a few gigabytes of FP16 memory.
@@
IV-B MCoT and M-ICL: Foundations of Multimodal Reasoning
MCoT reasoning extends the traditional CoT paradigm, originally developed for text-based LLMs, to multimodal settings by incorporating diverse data modalities, including images, videos, audio, speech, 3D data, tables, and charts. At its core, MCoT decomposes complex multimodal tasks into a sequence of intermediate reasoning steps, called thoughts, that together form a structured rationale before producing the final answer. Formally, given a multimodal input composed of language tokens $\mathcal{T}$ and non-linguistic modalities $\mathcal{M}$, MCoT can be categorized into two scenarios based on the nature of the rationale $\mathcal{R}$. In Scenario 1, the model generates text-only rationales to reason over multimodal inputs and outputs. In Scenario 2, the rationale itself incorporates multimodal elements, either as purely non-textual signals or as a hybrid of textual and multimodal content. By explicitly modeling the step-by-step progression from perception to cognition, this structured reasoning paradigm enhances interpretability, reduces hallucination, and improves performance on complex multimodal tasks.
The development of MCoT is closely linked to recent advances in multimodal LLMs such as GPT-4V, Gemini, and Claude3, which combine multimodal encoders with autoregressive decoders to process and generate cross-modal content. Despite this progress, several challenges remain, including the computational sustainability of long reasoning chains, error accumulation and propagation across intermediate steps, modality imbalance, hallucination mitigation, and the need for explainable and ethically aligned reasoning frameworks. Future research directions emphasize hybrid neuro-symbolic integration, adaptive reasoning in dynamic environments, and interdisciplinary insights from cognitive science, to advance MCoT toward more general and robust multimodal AGI [113].
M-ICL is another important emergent capability of large multimodal models. By presenting a small number of multimodal examples directly in the prompt, M-ICL enables models to solve novel tasks by analogy, significantly reducing reliance on large labeled datasets. Its training-free nature allows seamless deployment across a wide range of inference-time scenarios. Formally, M-ICL is defined as an inference-time adaptation mechanism for Large Multimodal Models (LMMs), in which the model processes a context $\mathcal{C}$ consisting of $\mathcal{N}$ multimodal demonstration tuples $(I_{i},T_{i},R_{i})$. Here, $I_{i}$ denotes an image, $T_{i}$ represents optional associated text (such as a question or instruction), and $R_{i}$ is the corresponding target response. This context is accompanied by a query $Q=(I_{Q},T_{Q})$, and the model generates an output $O=\text{LMM}(C,Q)$ under a causal language modeling objective, without any updates of gradient-based parameters.
@@ IV-B MCoT and M-ICL: Foundations of Multimodal Reasoning
IV-C MLLM-Assisted Human–Swarm Interaction and Autonomy
@@
Recent advances in MLLMs have significantly expanded the capabilities of Human–Swarm Interaction (HSI) and autonomous UAV systems. Eunmi et al.[33] highlight this shift by introducing SwarmChat, a context-aware interaction framework in which LLM-assisted modules, context generation, intent recognition, task planning, and modality selection collaborate to interpret user commands and adapt them to real-time swarm states. Their results demonstrate that natural-language, voice, and teleoperation interfaces can substantially reduce cognitive load. However, the system’s rule-based intent recognition and limited experimental scope indicate room for more scalable, learning-based solutions. Building on the idea that LLMs can mediate complex swarm behaviors, Liu et al.[77] present a multimodal GPT-4–based approach for UAV formation control using both visual and textual inputs. Their results show that MLLMs can extract commands from human instructions, interpret environmental cues from onboard imagery, and plan formations with over 80% success. This work validates the feasibility of end-to-end multimodal reasoning for swarm perception, coordination, and adaptive task execution, complementing the interaction-focused perspective introduced by SwarmChat. Extending multimodal reasoning into broader autonomous decision-making, Ping et al.[93] explore how MLLMs can enhance multi-agent perception, navigation, and coordination within UAV swarms. Through a forest-firefighting case study, they demonstrate that an MLLM (Doubao 1.5 Vision Pro) can perform natural-language task planning, dynamic fire assessment, and coordinated execution across multiple UAVs. Their discussion of challenges, such as hallucination, inference latency, and communication bottlenecks, further contextualizes the limitations identified in earlier work and underscores the need for more efficient, robust multimodal pipelines.
Table X: Multimodal LLM applications in UAV swarm operations, comparing frameworks for human-swarm interaction, autonomous formation, threat detection, natural language control, and emergency monitoring across diverse sensory and linguistic inputs.
@@
In addition to these operational advances, Xiao et al. [122] examine the role of MLLMs in secure and resilient swarm operations and propose a multimodal intrusion-detection architecture that integrates cyber and physical data streams. Their hierarchical zero-trust framework distributes different model sizes across edge, fog, and command layers, optimizing inference latency and computational cost while enhancing detection robustness. This security-focused perspective underscores the need for trustworthy MLLM integration as swarms become increasingly autonomous and data-intensive. From a control interface standpoint, Balaji et al. [7] introduce PromptPilot, which uses natural language instructions to control UAVs without requiring domain expertise. By combining a lightweight intent classifier with MLLMs such as Gemini or LLaVA, along with YOLOv10 and DeepSORT for vision-aware tasks, PromptPilot generates executable control code and demonstrates reliable autonomous performance. This work aligns with SwarmChat and Liu et al. in showing that intuitive human–UAV interaction can be achieved through MLLMs, while emphasizing practical deployment and real-world testing. Finally, Yang et al.[127] focus on rapid multimodal situational reporting, integrating UAV imagery, ground sensor data, and meteorological information to generate comprehensive forest-fire intelligence. Their system fuses Grounding DINO, YOLOv8-based models, and the DeepSeek LLM to identify critical environmental features and compute safe evacuation routes via enhanced A. This application illustrates how multimodal perception and language-based reasoning can support high-stakes environmental decision-making, complementing the swarm-control and interaction frameworks previously discussed. Table X provides a summary of these works.
Overall, SwarmChat integrates modules for context generation, intent recognition, task planning, and modality selection, enabling UAV swarms to interpret user commands and adapt to real-time operational states. Experiments show that natural language, voice, and teleoperation interfaces can substantially reduce cognitive load. However, current rule-based intent recognition and limited testing highlight the need for more scalable, learning-based solutions. Building on this approach, PromptPilot enables UAV control through natural language instructions without requiring domain expertise. By combining a lightweight intent classifier with MLLMs such as Gemini or LLaVA and vision-aware models such as YOLOv10 and DeepSORT, PromptPilot translates user instructions into executable control code and demonstrates reliable autonomous performance. Together, these systems illustrate the potential of MLLMs to mediate complex swarm behaviors, enhance human-machine interaction, and improve the efficiency, scalability, and resilience of UAV operations.
IV-D MLLM-Assisted UAV Navigation and Collaboration

### v2

@@ III-D Multi-LLM Architectures for UAV Communications
Multi-LLM systems are architectures in which multiple LLMs collaborate, either in parallel or sequentially, to solve complex tasks. Unlike single-model approaches, multi-LLM systems leverage the complementary strengths of individual models, enabling specialization, scalability, robustness, and effective task decomposition. Performance is further improved through cross-verification, which helps detect errors, reduce bias, and mitigate hallucinations. Collaboration among LLMs can follow cooperative, competitive, or ensemble paradigms, while system architectures may be centralized or decentralized, depending on fault-tolerance and task-allocation requirements. Additionally, multi-LLM frameworks can enhance privacy by routing sensitive queries to trusted models while validating outputs across agents.
@@
Extending multi-LLM coordination to interconnected aerial networks, Yan et al. [108] propose a hierarchical control framework for hybrid terrestrial–non-terrestrial systems incorporating High-Altitude Platform Stations (HAPS). Specifically, a meta-controller LLM deployed on the HAPS manages global resource allocation and load balancing, while onboard LLMs handle real-time motion planning and communication decisions. Experimental results show that this collaborative architecture achieves higher system rewards, lower operational costs, and lower collision rates than DRL-assisted and hybrid baselines. Fig. 6 depicts Multi-LLM architectures for UAV communications, and Table VII provides a summary of the works.
@@
Table VII: Hierarchical and multi-LLM architectures for coordinated UAV and aerial network operations, comparing distributed, swarm-based, and layered LLM frameworks in inspection, security, and hybrid terrestrial–non-terrestrial network applications.
@@ III-E Key Findings and Insights
Based on the above subsections, here we discuss the key findings and insights.
III-E1 LLM-assisted Spectrum Management
@@
Recent frameworks for LLM-assisted green-spectrum management in 6G Non-Terrestrial Networks (NTNs) integrate LLMs into a three-layer hierarchical architecture comprising data collection, intelligent LLM processing, and feedback-enabled execution. This design enables dynamic and adaptive spectrum allocation across satellites, high-altitude platforms, UAVs, and terrestrial networks. A key innovation is the joint optimization of spectral efficiency and energy consumption, supported by mechanisms including real-time interference analysis, predictive energy modeling, and an explicit energy–spectrum trade-off function. Simulation results show that the LLM-assisted approach significantly outperforms conventional methods, supporting up to 175 active transmitters – a 167% increase in capacity – while maintaining stable energy consumption and tighter interference control under varying threshold conditions. These findings advance intelligent and sustainable spectrum management for future 6G NTNs and underscore the transformative potential of LLMs in enabling green, scalable, high-performance global communications [83].
III-E2 LLM-assisted Near-Field Communications
@@
The integration of LLMs into near-field communications for the Low-Altitude Economy (LAE) brings transformative capabilities that address the unique challenges of UAV-centric wireless networks. By analyzing high-dimensional channel-state information in real time, LLMs can intelligently classify users as near-field or far-field without requiring precise distance estimation, thereby reducing computational overhead. They also enhance multi-user precoding by learning to predict beamforming and power allocation parameters directly from channel data, enabling efficient real-time adaptation to dynamic UAV mobility. Leveraging strong feature-extraction and pattern-recognition capabilities, LLMs can process the massive amounts of antenna and user data generated by extremely large-scale Multiple-Input Multiple-Output (MIMO) systems, thereby enabling performance to scale effectively as network density increases. In addition, their adaptability and few-shot learning abilities enable generalization across diverse operating conditions, such as varying UAV densities or interference levels, without extensive retraining. Finally, LLMs enable unified multi-task learning by jointly handling user classification, precoding, and interference management within a single model, thereby significantly simplifying the overall system design [106].
III-E3 LLM-assisted ISAC Optimization
@@
III-E4 High-Speed Connectivity for LLM-assisted UAVs
@@
III-E5 Assessing LLM Reasoning and Safety
Field deployments of integrated LLM-enabled UAV systems confirm significant improvements in operational adaptability and effectiveness across diverse applications, including emergency response, environmental monitoring, and infrastructure inspection. At the same time, these evaluations reveal persistent constraints in bandwidth availability, energy consumption, multimodal robustness, and reasoning reliability. These observations highlight the necessity of holistic co-design across communication, computation, and control, as well as the importance of standardized evaluation frameworks, such as UAVBench and UAVThreatBench, to objectively assess reasoning performance and cybersecurity readiness for safe and scalable deployment of intelligent UAV communication systems. The benchmark introduced in [32] addresses this gap by providing 50,000 validated UAV scenarios and an additional 50,000 multiple-choice questions ($UAVBench_{MCQ}$) spanning 10 distinct reasoning styles. This comprehensive design enables systematic evaluation of a wide range of cognitive capabilities.
@@ IV-A Building and Fine-Tuning Multimodal LLMs
A key challenge lies in aligning heterogeneous modalities for grounded reasoning. Contemporary MLLMs typically rely on modality-specific encoders coupled with a pretrained LLM, thereby enabling language modules to selectively attend to visual features. This flexible architecture serves as a common backbone for UAV perception and reasoning tasks, but its effectiveness in aerial settings depends strongly on downstream adaptation.
Following pretraining, models are fine-tuned on UAV-specific tasks such as aerial image captioning, visual question answering, and object recognition. Dedicated datasets, including AeroCaps [10] and HRVQA-VL [35], support fine-grained scene understanding and multi-hop reasoning over UAV imagery. Empirical evidence shows that task-aware fine-tuning is critical: compact MLLMs (e.g., 2B parameters) adapted with UAV data can outperform significantly larger generic models while requiring only a few gigabytes of FP16 memory.
@@
IV-B MLLM-Assisted Human–Swarm Interaction and Autonomy
@@
Recent advances in MLLMs have significantly expanded the capabilities of Human–Swarm Interaction (HSI) and autonomous UAV systems. Eunmi et al.[31] highlight this shift by introducing SwarmChat, a context-aware interaction framework in which LLM-assisted modules, context generation, intent recognition, task planning, and modality selection collaborate to interpret user commands and adapt them to real-time swarm states. Their results demonstrate that natural-language, voice, and teleoperation interfaces can substantially reduce cognitive load. However, the system’s rule-based intent recognition and limited experimental scope indicate room for more scalable, learning-based solutions. Building on the idea that LLMs can mediate complex swarm behaviors, Liu et al.[69] present a multimodal GPT-4–based approach for UAV formation control using both visual and textual inputs. Their results show that MLLMs can extract commands from human instructions, interpret environmental cues from onboard imagery, and plan formations with over 80% success. This work validates the feasibility of end-to-end multimodal reasoning for swarm perception, coordination, and adaptive task execution, complementing the interaction-focused perspective introduced by SwarmChat. Extending multimodal reasoning into broader autonomous decision-making, Ping et al.[80] explore how MLLMs can enhance multi-agent perception, navigation, and coordination within UAV swarms. Through a forest-firefighting case study, they demonstrate that an MLLM (Doubao 1.5 Vision Pro) can perform natural-language task planning, dynamic fire assessment, and coordinated execution across multiple UAVs. Their discussion of challenges, including hallucinations, inference latency, and communication bottlenecks, further contextualizes the limitations identified in earlier work and underscores the need for more efficient, robust multimodal pipelines.
Table VIII: Multimodal LLM applications in UAV swarm operations, comparing frameworks for human-swarm interaction, autonomous formation, threat detection, natural language control, and emergency monitoring across diverse sensory and linguistic inputs.
@@
In addition to these operational advances, Xiao et al. [104] examine the role of MLLMs in secure and resilient swarm operations and propose a multimodal intrusion-detection architecture that integrates cyber and physical data streams. Their hierarchical zero-trust framework distributes models of varying sizes across edge, fog, and command layers, optimizing inference latency and computational cost while enhancing detection robustness. This security-focused perspective underscores the need for trustworthy MLLM integration as swarms become increasingly autonomous and data-intensive. From a control interface standpoint, Balaji et al. [9] introduce PromptPilot, which uses natural language instructions to control UAVs without requiring domain expertise. By combining a lightweight intent classifier with MLLMs such as Gemini or LLaVA, along with YOLOv10 and DeepSORT for vision-aware tasks, PromptPilot generates executable control code and demonstrates reliable autonomous performance. This work aligns with SwarmChat and Liu et al. in showing that intuitive human–UAV interaction can be achieved through MLLMs, while emphasizing practical deployment and real-world testing. Finally, Yang et al.[109] focus on rapid multimodal situational reporting, integrating UAV imagery, ground sensor data, and meteorological information to generate comprehensive forest-fire intelligence. Their system fuses Grounding DINO, YOLOv8-based models, and the DeepSeek LLM to identify critical environmental features and compute safe evacuation routes via enhanced A. This application illustrates how multimodal perception and language-based reasoning can support high-stakes environmental decision-making, complementing the swarm-control and interaction frameworks previously discussed. Table VIII provides a summary of these works.
Overall, SwarmChat integrates modules for context generation, intent recognition, task planning, and modality selection, enabling UAV swarms to interpret user commands and adapt to real-time operational states. Experiments show that natural language, voice, and teleoperation interfaces can substantially reduce cognitive load. However, current rule-based intent recognition and limited testing highlight the need for more scalable, learning-based solutions. Building on this approach, PromptPilot enables UAV control via natural-language instructions without requiring domain expertise. By combining a lightweight intent classifier with MLLMs such as Gemini or LLaVA and vision-aware models such as YOLOv10 and DeepSORT, PromptPilot translates user instructions into executable control code and demonstrates reliable autonomous performance. Together, these systems illustrate the potential of MLLMs to mediate complex swarm behaviors, enhance human-machine interaction, and improve the efficiency, scalability, and resilience of UAV operations.
IV-C MLLM-Assisted UAV Navigation and Collaboration